# FedVSR: Towards Model-Agnostic Federated Learning in Video Super-Resolution

### v2

@@ Abstract
Video super-resolution aims to enhance low-resolution videos by leveraging both spatial and temporal information. While deep learning has led to impressive progress, it typically requires centralized data, which raises privacy concerns. Federated learning offers a privacy-friendly solution, but general FL frameworks often struggle with low-level vision tasks, resulting in blurry, low-quality outputs. To address this, we introduce FedVSR, the first FL framework specifically designed for VSR. It is model-agnostic and stateless, and introduces a lightweight loss function based on the DWT to better preserve high-frequency details during local training. Additionally, a loss-aware aggregation strategy combines both DWT-based and task-specific losses to guide global updates effectively. Extensive experiments across multiple VSR models and datasets demonstrate that FedVSR consistently outperforms existing FL methods, achieving up to 0.82 dB higher PSNR, 0.0327 higher SSIM, and 0.0251 lower LPIPS. These results underscore FedVSR’s ability to bridge the gap between privacy and performance, setting a new benchmark for federated learning in low-level vision tasks. The code is available at: https://github.com/alimd94/FedVSR
@@ 1 Introduction
Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 30.41, 0.8600, 0.1777 Refer to caption (d) FedProx[27] 30.07, 0.8531, 0.1851
@@
Federated learning (FL) provides a principled remedy to these concerns [36]. In many VSR scenarios, such as smartphones, surveillance, health-related domains, and remote edge sensors, videos are captured on devices with limited bandwidth and strict privacy constraints. Uploading hours of footage can consume tens of gigabytes of storage; for instance, even a short 10-second 8K video at 30 fps in raw format can exceed 10 GB in size [2]. Additionally, uploading such data may violate policies and laws such as GDPR or HIPAA [19]. Unlike conventional distributed methods that rely on frequent synchronization, FL enables selected participants to perform multiple local updates before their models are aggregated into a global one. This approach minimizes communication overhead and accommodates the limited computational resources of edge devices. These combined benefits make FL a logical and effective foundation for VSR systems that face such concerns.
@@
Despite significant advancements in FL methods, most research has focused on high-level or downstream visual tasks such as image classification [21, 27, 51], facial recognition [41, 34], and segmentation [37, 34]. However, their application in super-resolution (SR) remains largely unexplored. Existing studies [39, 56] found that directly applying existing general FL techniques to SR tasks results in poor-quality reconstructions with blurred textures, mainly due to their inability to effectively capture rich low-level information such as texture, edges and motion. Motivated by these limitations, we propose FedVSR, a novel model-agnostic framework for VSR under FL. FedVSR treats VSR models as black boxes, allowing it to work with a variety of architectures without requiring architectural changes or access to intermediate layers. To improve performance under FL constraints, we introduce a lightweight loss term based on the Discrete Wavelet Transform (DWT), which emphasizes high-frequency details that are often lost during decentralized training. While high-frequency preservation is valuable for all SR tasks, its importance is amplified in VSR due to the added complexity of maintaining temporal coherence across frames. In federated settings, data heterogeneity and limited communication intensify this challenge. FedVSR addresses these issues without maintaining state between rounds, making it both scalable and practical for privacy-sensitive, resource-constrained environments. In summary, the main contributions of this paper are as follows:
  • We design a frequency-aware loss using DWT to sharpen fine details during client updates, and introduce a complementary aggregation method to weight model contributions more effectively. To the best of our knowledge, this is the first algorithm specifically designed to address VSR in the FL setting, effectively bridging the gap between privacy preservation and reconstruction quality.
@@
  • Extensive experiments validate FedVSR’s effectiveness, outperforming existing approaches and paving the way for privacy-preserving low-level vision in the FL setting.
@@ 2.1 Federated Learning
Federated learning (FL) is a decentralized learning paradigm that enables multiple clients to collaboratively train models while preserving data privacy. It has been widely applied across diverse domains, including healthcare, finance, and IoT, where data security and regulatory constraints are critical [3, 7, 62]. A foundational approach in this domain is FedAvg [36], which follows a four-step iterative process: (1) the server distributes a global model to all participating clients, (2) each client trains its local model using its own data, (3) the locally updated models are sent back to the central server, and (4) the server aggregates these updates to refine the global model for the next training round.
@@
To enhance the performance of federated learning, multiple methods focus on optimizing local updates. FedProx [27] incorporates a proximal term in the local objective to ensure consistency between local and global models. SCAFFOLD [21] introduces control variates for both clients and the server. FedNova [51] employs normalized averaging to address objective inconsistencies. More recently, MOON [25] applies contrastive learning to assess similarity among different model representations.
2.2 Video Super Resolution
Early VSR methods that applied SISR techniques frame-by-frame neglected the temporal context, resulting in temporal inconsistencies, loss of detail, and undesired smoothing in the high-resolution output [33]. VSR leverages multiple temporally correlated low-resolution frames within a video sequence to generate high-resolution output. Unlike SISR, VSR must account for both spatial and temporal dependencies across frames, making it a highly complex, non-linear, multi-dimensional problem.
2.3 Federated Learning on Super Resolution Task
@@
For low-level vision tasks, FedMRI [15] introduces a federated learning framework for MRI reconstruction with a shared encoder for generalized features and client-specific decoders for personalization. FedPR [16] improves image quality using pre-trained models and prompt learning with minimal trainable parameters. p²FedSR [55] enhances privacy-preserving SR by integrating client-specific degradation priors and a selective model aggregation strategy for personalized FL. FedSR [56] employs a detail-assisted contrastive loss for feature alignment at the client level and a hierarchical aggregation policy at the server level with historical global model, optimizing layer-wise model integration.
3 Preliminaries
3.1 VSR Problem Definition
@@
VSR enhances video resolution by using multiple low-resolution frames, leveraging temporal information to exploit inter-frame dependencies, unlike SISR.
Problem Formulation
In addition to the widely used RGB color space, other representations such as the YUV and YCbCr color spaces are frequently employed in VSR applications. Let $F_{i}\in\mathbb{R}^{H\times W\times 3}$ denote the $i$-th frame in a given LR video sequence, while $\hat{F}_{i}\in\mathbb{R}^{sH\times sW\times 3}$ represents its corresponding high-resolution (HR) counterpart, where $s$ is the scaling factor (e.g., $s=2,4,8$). Additionally, we define a set of $12N+$ 1 HR frames $\{\hat{I}_{j}\}_{j=i-N}^{i+N}$ centered around $\hat{I}_{i}$, where $N$ represents the temporal radius. The degradation process that generates an LR video sequence from its HR counterpart can be mathematically formulated as:
$I_{i}=\phi(\hat{I}_{i},\{\hat{I}_{j}\}_{j=i-N}^{i+N};\theta_{\alpha})$ (1)
where $\phi(\cdot;\cdot)$ is the degradation function parameterized by $\theta_{\alpha}$, incorporating various degradation factors such as noise, motion blur, and down-sampling effects.
A more detailed expression for the degradation process can be written as:
$I_{j}=DBE_{i\to j}(\hat{I}_{i})+n_{j}$ (2)
where $D$ represents the down-sampling operation, $B$ denotes the blurring effect, $E_{i\to j}$ is a warping function that accounts for motion between frames $\hat{I}_{i}$ and $\hat{I}_{j}$, and $n_{j}$ represents the image noise. The goal of a VSR algorithm is to reconstruct an HR video sequence from degraded LR frames in a manner that closely resembles the ground truth video. This makes this task an ill-posed problem because there exist infinitely many possible HR images $\hat{I}_{i}$ that can produce the same LR image $I_{i}$ after degrading, meaning the degradation operator is highly non-injective [49]. While some of the techniques used in ISIR can be applied to individual frames, VSR must also consider inter-frame dependencies to ensure motion coherence across the sequence. The inverse process of degradation, which represents the super-resolution task, can be expressed as:
$\tilde{I}_{i}=\phi^{-1}(I_{i},\{I_{j}\}_{j=i-N}^{i+N};\theta_{\beta})$ (3)
where $\tilde{I}_{i}$ is the estimated HR frame corresponding to the ground truth $\hat{I}_{i}$, and $\theta_{\beta}$ represents the model parameters involved in the reconstruction process [33].
3.2 FL Problem Definition
@@
FL enables multiple clients to train a shared model collaboratively without sharing raw data, enhancing privacy and reducing communication overhead through local computations and periodic model aggregation.
Problem Formulation
FL aims to minimize the following objective function across participating clients:
$\min_{\mathbf{w}}F(\mathbf{w})=\sum_{k=1}^{K}p_{k}F_{k}(\mathbf{w})$ (4)
where:
  • $K$ is the total number of clients,
  • $p_{k}$ is the relative weight of client $k$ (typically based on its dataset size),
  • $F_{k}(\mathbf{w})$ is the local loss function for client $k$,
  • $\mathbf{w}$ represents the global model parameters.
Each client updates the global model locally:
$\mathbf{w}_{k}^{t+1}=\mathbf{w}^{t}-\eta\nabla F_{k}(\mathbf{w}^{t})$ (5)
where $\eta$ is the learning rate. The updated models are aggregated by the central server using a weighted averaging scheme:
$\mathbf{w}^{t+1}=\sum_{k=1}^{K}p_{k}\mathbf{w}_{k}^{t+1}$ (6)
Not all clients participate equally in every training round due to data heterogeneity, resource constraints, network instability, or other limitations [59]. Client averaging based on their contributions stabilizes training, reduces variance, and improves generalization [6].
However, weighting clients solely by dataset size, as in traditional FedAvg[36], does not fully capture their impact on model performance [44]. A client’s importance extends beyond data quantity and is also influenced by the quality of its model updates and the loss it generates during training. Based on [46], we implement an approach where clients with lower loss are typically better trained and can be assigned higher weights during averaging. This ensures improved overall performance by leveraging the robustness properties analyzed in the paper. It is worth noting that some works instead prioritize clients with higher loss to increase fairness [26]. This approach, which contrasts with a focus on final performance, is more common in the context of personalized federated learning.
By incorporating loss-based importance into the aggregation process, federated learning can achieve more reliable and efficient training. This adaptive weighting ensures that client updates more accurately reflect their significance in improving the global model, leading to a more stable and equitable learning process [17].
4 Methodology
VSR remains a challenging problem due to its ill-posed nature. Severe degradations amplify artifacts, making it difficult to balance detail enhancement and noise suppression. To improve generalizability, diverse training data and large batch sizes are needed, but this significantly increases computational costs [8]. FL adds further complexity to VSR due to real-world constraints [14] and the nature of the task. In heterogeneous environments, differences in client data distributions lead to biased updates, poor generalization [5], and inefficient model merging. Gradient mismatches [53] and parameter inconsistencies slow convergence [10]. Most FL algorithms are designed for image classification which can be considered as a well-posed problem [18]. However, VSR is an ill-posed task, requiring a further fundamental assessment of existing methods to ensure effective adaptation. For instance, FedProx assumes a certain degree of similarity across client data distributions, but in real-world applications, these distributions can vary significantly. When this assumption fails, models experience drift, leading to inconsistent local updates and poor generalization [13]. Similarly, SCAFFOLD assumes unbiased client updates, yet in large-scale deployments, heterogeneous data causes significant divergence, slowing convergence. Moreover, its need for additional state information per client makes it impractical for large-scale FL systems due to communication and storage constraints [10]. Meanwhile, Other methods such as FedDisco[58] and MOON rely on class label distributions or logits, making them inherently task-specific rather than generalizable.This work proposes a solution to enhance the overall objective of VSR models while improving aggregation to mitigate these challenges.
4.1 Enhancing Loss Function for Federated VSR
@@
4.1.1 DWT for High-Frequency Preservation
@@
One common failure in training VSR models is producing overly smooth outputs, meaning the trained model fails to capture high-frequency information. As a result, details and textures are missing from the output. Since the models we used are among the top-performing ones in a centralized setting, this suggests that the failure to restore high-frequency details is specifically related to training under the FL paradigm. To alleviate this, we leverage the Discrete Wavelet Transform (DWT), which has proven effective in super-resolution tasks [20, 61].
DWT is a multi-resolution signal decomposition technique that represents a signal in terms of approximation and detail coefficients at different scales. For a 2D signal (such as an image or video frame), DWT decomposes the signal into sub-bands of different frequencies and orientations such as:
@@
  • Approximation Coefficients $F_{i}^{LL}$ (low-pass filtered in both dimensions).
  • Horizontal Detail Coefficients $F_{i}^{LH}$ (high-pass in vertical, low-pass in horizontal).
  • Vertical Detail Coefficients $F_{i}^{HL}$ (low-pass in vertical, high-pass in horizontal).
  • Diagonal Detail Coefficients $F_{i}^{HH}$ (high-pass in both dimensions).
One of the simplest wavelets used in DWT is Haar wavelet, which consists of a step function. It is computationally efficient and commonly used in image and video processing tasks. In this method each frame’s pixel values are convolved with:
  • Low-pass filter $h=\frac{1}{\sqrt{2}}[1,1]$ to extract approximation coefficients.
  • High-pass filter $g=\frac{1}{\sqrt{2}}[1,-1]$ to extract detail coefficients.
This process is repeated independently for each frame in the sequence.
The coefficient in those four sub-bands are calculated as:
$F_{i}^{LL}=(F_{i}*h^{T})*h,\quad F_{i}^{LH}=(F_{i}*h^{T})*g$ (7)
$F_{i}^{HL}=(F_{i}*g^{T})*h,\quad F_{i}^{HH}=(F_{i}*g^{T})*g$ (8)
The low-frequency component $F_{i}^{LL}$, contains the content and color information of the input frame, while the three high-frequency sub-bands capture detailed information about global structures and textures [43]. Therefore, for our loss helper term, we consider only the high-frequency sub-bands and ignore the low-frequency ones.
The loss is then computed pixel-wise in each sub-band between the predicted $\bar{F}$ and ground truth $\hat{F}$ frame at time step $i$:
$L_{i}^{LH}=\bar{F}_{i}^{LH}-\hat{F}_{i}^{LH}$ (9)
$L_{i}^{HL}=\bar{F}_{i}^{HL}-\hat{F}_{i}^{HL}$ (10)
$L_{i}^{HH}=\bar{F}_{i}^{HH}-\hat{F}_{i}^{HH}$ (11)
Then with the Charbonnier loss [9] we calculate the helper term $\mathcal{L}_{\text{DWT}}$ :
@@
$\mathcal{L}_{\text{HiFr}}=\sum_{i}\sum_{s\in\{LH,HL,HH\}}\sqrt{{L_{i}^{s}}^{2}+\epsilon^{2}}$ (12)
Now, we define the overall loss for the model with two weighted terms, one for the generic VSR loss $\mathcal{L}_{\text{VSR}}$ and one for the DWT-based high-frequency loss $\mathcal{L}_{\text{HiFr}}$ as follows:

### v3

@@ Abstract
Video super-resolution (VSR) aims to enhance low-resolution videos by leveraging both spatial and temporal information. While deep learning has led to impressive progress, it typically requires centralized data, which raises privacy concerns. Federated learning (FL) offers a privacy-friendly solution, but general FL frameworks often struggle with low-level vision tasks, resulting in blurry, low-quality outputs. To address this, we introduce FedVSR, the first FL framework specifically designed for VSR. It is model-agnostic and stateless, and introduces a lightweight loss function based on the Discrete Wavelet Transform (DWT) to better preserve high-frequency details during local training. Additionally, a loss-aware aggregation strategy combines both DWT-based and task-specific losses to guide global updates effectively. Extensive experiments across multiple VSR models and datasets show that FedVSR not only improves perceptual video quality (up to +0.89 dB PSNR, +0.0370 SSIM, 0.0347-0.0347 LPIPS and 4.98 VMAF) but also achieves these gains with close to zero computation and communication overhead compared to its rivals. These results demonstrate FedVSR’s potential to bridge the gap between privacy, efficiency, and perceptual quality, setting a new benchmark for federated learning in low-level vision tasks. The code is available at: https://github.com/alimd94/FedVSR
@@ 1 Introduction
Refer to caption (b) FedAvg[30] 30.41, 0.8600, 0.1777 Refer to caption (c) FedMedian [52] 28.89, 0.8224, 0.2106 Refer to caption (d) FedProx[21] 30.07, 0.8531, 0.1851
@@
Federated learning (FL) provides a principled remedy to these concerns [30]. In numerous Video Super-Resolution (VSR) scenarios, such as remote patient monitoring, personal phone footage analysis, and confidential surveillance, video data captured on edge devices is inherently privacy-sensitive. Training on this data centrally is often prohibited by stringent regulations such as GDPR or CCPA [25], which forbid uploading and aggregating raw footage containing personally identifiable information. Privacy is the core motivator for FL in VSR. This decentralization also solves the massive logistical challenge of video data size; for instance, attempting to upload even a short 10-second 8K video for training can easily exceed 10 GB [2]. Unlike conventional distributed methods that rely on frequent synchronization, FL enables selected participants to perform multiple local updates before their models are aggregated into a global one. This approach minimizes communication overhead and accommodates the limited computational resources of edge devices. These combined benefits make FL a logical and effective foundation for VSR systems that face such concerns.
@@
Despite significant advancements in FL methods, most research has focused on high-level or downstream visual tasks such as image classification [18, 21, 46], facial recognition [29], and segmentation [31, 29]. However, their application in super-resolution (SR) remains largely unexplored. Existing studies [34, 50] found that directly applying general FL techniques to low-level vision tasks like SR results in poor-quality reconstructions with blurred textures. This failure is compounded in VSR, where data heterogeneity (diverse video content and quality) and communication constraints severely degrade the model’s ability to capture and synchronize rich spatial and temporal high-frequency details across frames, a prerequisite for high-fidelity video output. Motivated by these limitations, we propose FedVSR, a novel framework for VSR under FL. FedVSR is model-agnostic, meaning it treats underlying VSR models as black boxes and works by modifying only the local training loss and the server aggregation strategy. This enables compatibility with diverse VSR architectures without requiring internal model modifications or access to intermediate layers, which is crucial for scalability and flexibility in a real-world federated deployment. The framework consists of two key components. The first component is a novel, frequency-aware training phase enhanced by a lightweight 3D Discrete Wavelet Transform (DWT)-based loss. By operating in the frequency domain, this loss forces clients to prioritize the spatial and temporal high-frequency details essential for VSR that are typically lost due to FL’s limited communication budget. Crucially, we demonstrate that this loss is specific to the FL environment, as its direct application in centralized VSR leads to performance degradation (see Section 5.4.1). The second is an innovative adaptive loss aggregation mechanism that explicitly uses client training quality to mitigate the detrimental effects of data heterogeneity and noisy clients. While high-frequency preservation is important for all SR tasks, it is especially crucial in VSR to maintain temporal coherence across frames. Federated settings further amplify this challenge due to data heterogeneity and limited communication. By combining its two components, FedVSR enhances training and aggregation robustness while remaining scalable, state-free, and practical for privacy-sensitive, resource-constrained environments. In summary, the main contributions of this paper are as follows:
  • We design a two-pronged solution to the Federated VSR challenge: (1) a novel, frequency-aware loss using 3D DWT to preserve critical spatial-temporal high-frequency details during decentralized training, and (2) an adaptive loss-aware aggregation method that dynamically adjusts client weights to combat data heterogeneity. This combination represents the first holistic framework designed to bridge the gap between VSR quality and FL privacy.
@@
  • Extensive experiments validate FedVSR’s effectiveness and robustness among existing approaches and paving the way for privacy-preserving low-level vision in the FL setting.
Figure 2: Overview of the proposed FedVSR framework. Each client computes a model-agnostic VSR update augmented with a DWT-based high-frequency loss. Clients also track the average local loss , which is used for loss-aware weighted aggregation at the server. The global model is iteratively refined while maintaining model-agnostic and stateless properties.
@@ 2 Related Works
2.1 Video Super-Resolution
Early VSR methods that applied SISR techniques frame-by-frame neglected the temporal context, resulting in temporal inconsistencies, loss of detail, and undesired smoothing in the high resolution output [28]. VSR leverages multiple temporally correlated low-resolution frames within a video sequence to generate high-resolution output. Unlike SISR, VSR must account for both spatial and temporal dependencies across frames, making it a highly complex, non-linear, multi-dimensional problem.
@@ 2.2 Federated Learning
Federated learning (FL) is a decentralized paradigm that enables multiple clients to collaboratively train models while preserving data privacy. It has been widely applied in domains such as healthcare, finance, and the Internet of Things (IoT), where data security and regulatory compliance are critical [3, 7, 57]. A foundational approach is FedAvg [30], which follows a four-step iterative process: (1) the server distributes a global model to all clients, (2) each client trains its local model on its own data, (3) locally updated models are sent back to the server, and (4) the server aggregates these updates to refine the global model for the next round.
@@
To improve FL performance, several methods focus on optimizing local updates. FedProx [21] adds a proximal term to maintain consistency between local and global models. SCAFFOLD [18] introduces control variates for both clients and server, while FedNova [46] employs normalized averaging to address objective inconsistencies. More recently, MOON [20] leverages contrastive learning to assess similarity among different model representations.
Although these methods enhance convergence and stability in classification tasks, their reliance on class-specific assumptions limits applicability to low-level, regression-based tasks such as VSR. For instance, MOON, FedDyn [1], FedDisco [51], and FedLAW [22] depend on logits, class-wise statistics, or server-side data proxies—assumptions that do not translate to VSR or SISR scenarios.
2.3 Federated Learning in Image Super Resolution
@@
Based on our review of the literature, FL has not yet been explored for VSR; the works discussed below are currently the only FL-based approaches applied to SISR. For Image-based low-level vision tasks, FedMRI [14] introduces a federated learning framework for MRI reconstruction using a shared encoder for generalized features and client-specific decoders for personalization. FedPR [15] leverages pre-trained models and prompt learning with minimal trainable parameters to improve image quality. p²FedSR [49] enhances personalized privacy-preserving SR through client-specific degradation priors and selective model aggregation, while FedSR [50] employs detail-assisted contrastive loss for feature alignment and a hierarchical aggregation policy for layer-wise model integration.
Although these approaches represent significant progress in FL for low-level vision, they are restricted to SISR rather than VSR. Moreover, methods such as FedSR, p²FedSR, and FedMRI are closely tied to specific model architectures (shallow——deep——upsampler), limiting applicability to a narrow range of SR frameworks. They also rely on stateful training: clients must retain auxiliary information such as degradation priors, feature statistics, or historical checkpoints, which introduces non-trivial storage, computation, and synchronization overhead. Such constraints hinder scalability and deployment, especially in heterogeneous or resource-limited environments.
In this paper, we propose FedVSR, a model-agnostic and stateless framework that allows any VSR model to be seamlessly incorporated without structural modifications or client-side state management. This design ensures easy extensibility, offering greater flexibility and scalability than prior work, and making FedVSR practical for real-world deployment where lightweight, generalizable solutions are essential.
3 Overview of FedVSR
FedVSR integrates Video Super-Resolution (VSR) with Federated Learning (FL), enabling collaborative enhancement of video resolution across distributed clients while preserving data privacy. This section provides a concise overview of the underlying VSR and FL problems and their formulations.
3.1 Video Super-Resolution (VSR)
@@
VSR aims to enhance video resolution by leveraging temporal information across multiple low-resolution (LR) frames, exploiting inter-frame dependencies that single-image super-resolution (SISR) methods cannot utilize:
Let $F_{i}\in\mathbb{R}^{H\times W\times 3}$ denote the $i$-th frame in an LR video sequence, and $\hat{F}_{i}\in\mathbb{R}^{sH\times sW\times 3}$ its high-resolution (HR) counterpart, where $s$ is the upscaling factor. Define a temporal window of $2N+1$ HR frames $\{\hat{I}_{j}\}_{j=i-N}^{i+N}$ centered around $\hat{I}_{i}$. The LR frames are generated via a degradation process:
$I_{i}=\phi(\hat{I}_{i},\{\hat{I}_{j}\}_{j=i-N}^{i+N};\theta_{\alpha}),$ (1)
where $\phi(\cdot;\cdot)$ models degradation factors such as blur, noise, and motion. A more explicit formulation is:
$I_{j}=DBE_{i\to j}(\hat{I}_{i})+n_{j},$ (2)
with $D$ as the down-sampling operator, $B$ the blur effect, $E_{i\to j}$ a warping function for motion compensation, and $n_{j}$ image noise. VSR seeks to reconstruct HR frames:
$\tilde{I}_{i}=\phi^{-1}(I_{i},\{I_{j}\}_{j=i-N}^{i+N};\theta_{\beta}),$ (3)
where $\tilde{I}_{i}$ approximates the ground truth $\hat{I}_{i}$ [44, 28]. This is an ill-posed problem due to the non-injective nature of the degradation function.
3.2 Federated Learning (FL)
@@
FL allows multiple clients to collaboratively train a shared model without exchanging raw data, reducing privacy risks and communication overhead. The global optimization problem is:
$\min_{\mathbf{w}}F(\mathbf{w})=\sum_{k=1}^{K}p_{k}F_{k}(\mathbf{w}),$ (4)
where $K$ is the number of clients, $p_{k}$ is the relative weight of client $k$, $F_{k}(\mathbf{w})$ the local loss, and $\mathbf{w}$ the global model parameters. Clients update locally via:
$\mathbf{w}_{k}^{t+1}=\mathbf{w}^{t}-\eta\nabla F_{k}(\mathbf{w}^{t}),$ (5)
and the server aggregates updates with weighted averaging:
$\mathbf{w}^{t+1}=\sum_{k=1}^{K}p_{k}\mathbf{w}_{k}^{t+1}.$ (6)
Client participation can vary due to data heterogeneity, resource limits, or network instability, and simple dataset-size weighting may not capture each client’s true impact [54, 5, 30, 38]. FedVSR incorporates hybrid adaptive loss-based weighting, where clients with lower local loss contribute more effectively to the global model, improving training stability and performance. See Section 4.2
3.3 FedVSR: Integrating VSR with FL
FedVSR combines these paradigms by distributing the VSR training process across multiple clients. Each client improves HR frame reconstruction locally while preserving its private video data. The server aggregates model updates using adaptive weights to optimize global performance. This approach maintains temporal coherence in videos, leverages multi-client data diversity, and preserve privacy to address the challenges of centralized VSR training.
4 Proposed FedVSR Framework
VSR remains a challenging problem due to its ill-posed nature. Severe degradations amplify artifacts, making it difficult to balance detail enhancement and noise suppression. To improve generalizability, diverse training data and large batch sizes are needed, but this significantly increases computational costs [8]. FL adds further complexity to VSR due to real-world constraints [13] and the nature of the task. In heterogeneous environments, differences in client data distributions lead to biased updates, poor generalization [4], and inefficient model merging. Gradient mismatches [47] and parameter inconsistencies slow convergence [10]. Our design is a two-pronged approach that enhances client training via a novel loss term and improves server robustness through adaptive aggregation. The overall architecture of FedVSR framework is illustrated in Fig. 2
4.1 High-Frequency Loss for Federated VSR
@@
4.1.1 3D DWT for High-Frequency Preservation
@@
One common failure in training VSR models is producing overly smooth outputs, meaning the trained model fails to capture high-frequency information. As a result, details and textures are missing from the output. Since the models we used are among the top-performing ones in a centralized setting, this suggests that the failure to restore high-frequency details is specifically related to training under the FL paradigm. To alleviate this, we leverage the 3D Discrete Wavelet Transform (DWT). While 2D DWT has shown effectiveness in SISR [17, 56], we apply the 3D variant to capture temporal high-frequency details across frames in addition to spatial ones. This is essential for maintaining temporal coherence in VSR output, a problem standard FL aggregation fails to address effectively.
DWT is a multi-resolution signal decomposition technique that represents a signal in terms of approximation and detail coefficients at different scales. For our implementation, we select the separable Haar wavelet. While more sophisticated wavelets offer better decorrelation, Haar’s inherent simplicity and use of short integer filters result in minimal computational overhead during the local update step. This is a critical design choice for FL, as it aligns with our goal of introducing a lightweight, non-intensive loss term that does not significantly burden the resource-constrained client devices. For a 3D signal (such as a volumetric image or a video clip), DWT decomposes the signal into eight sub-bands of different frequencies and orientations:
@@
  • Low-Frequency Approximation Coefficients $F_{i}^{LLL}$ (low-pass filtered along all three dimensions).
  • High-Frequency Coefficients with single high-pass filtering:
    • $F_{i}^{LLH}$ (low-pass depth and height, high-pass width),
    • $F_{i}^{LHL}$ (low-pass depth and width, high-pass height),
    • $F_{i}^{HLL}$ (low-pass height and width, high-pass depth).
  • High-Frequency Coefficients with two high-pass filterings:
    • $F_{i}^{LHH}$ (low-pass depth, high-pass height and width),
    • $F_{i}^{HLH}$ (low-pass height, high-pass depth and width),
    • $F_{i}^{HHL}$ (low-pass width, high-pass depth and height).
  • Full High-Frequency Coefficients $F_{i}^{HHH}$ (high-pass in all three dimensions).
Using the separable Haar wavelet, the 1D filters are defined as:
  • Low-pass filter $h=\frac{1}{\sqrt{2}}[1,1]$,
  • High-pass filter $g=\frac{1}{\sqrt{2}}[1,-1]$.
The coefficients in the eight sub-bands are computed by successive filtering along the three dimensions (depth, height, width). For example:
$F_{i}^{LLL}$ $=((F_{i}*h^{T})*h)*h$ (7)
$F_{i}^{LLH}$ $=((F_{i}*h^{T})*h)*g$ (8)
$F_{i}^{LHL}$ $=((F_{i}*h^{T})*g)*h$ (9)
$F_{i}^{LHH}$ $=((F_{i}*h^{T})*g)*g$ (10)
$F_{i}^{HLL}$ $=((F_{i}*g^{T})*h)*h$ (11)
$F_{i}^{HLH}$ $=((F_{i}*g^{T})*h)*g$ (12)
$F_{i}^{HHL}$ $=((F_{i}*g^{T})*g)*h$ (13)
$F_{i}^{HHH}$ $=((F_{i}*g^{T})*g)*g.$ (14)
Similar to the 2D case, the low-frequency component $F_{i}^{LLL}$ captures the global structure and coarse content of the input, while the seven high-frequency sub-bands capture edges, textures, and structural variations along different orientations [37].
For our loss design, we focus on the high-frequency sub-bands and ignore the low-frequency approximation. The high-frequency helper loss is then computed using the Charbonnier loss [9]:
@@
$\mathcal{L}_{\mathrm{HiFr}}(\bar{F},\hat{F})=\mathbb{E}\!\left[\sqrt{\big(H_{\text{HF}}(\bar{F})-H_{\text{HF}}(\hat{F})\big)^{2}+\epsilon^{2}}\right]$ (15)
where $H_{\text{HF}}(\cdot)$ denotes the concatenation of those seven high frequency sub-bands $\{LLH,LHL,LHH,HLL,HLH,HHL,HHH\}$ obtained by the 3D DWT from the predicted $\bar{F}$ and ground truth $\hat{F}$ frame , $\epsilon>0$ is the Charbonnier constant, and $\mathbb{E}[\cdot]$ indicates the mean over all elements (batch, channels, temporal, and spatial).
Finally, the total training objective combines the standard VSR loss with the DWT-based high-frequency loss: